# SWE-SQL: Illuminating LLM Pathways to Solve User SQL Issues in Real-World Applications

### v2

@@ Abstract
Resolution of complex SQL issues persists as a significant bottleneck in real-world database applications. Current Large Language Models (LLMs), while adept at text-to-SQL translation, have not been rigorously evaluated on the more challenging task of debugging on SQL issues. In order to address this gap, we introduce BIRD-CRITIC, a new SQL issue debugging benchmark comprising 530 carefully curated PostgreSQL tasks (BIRD-CRITIC-PG) and 570 multi-dialect tasks (BIRD-CRITIC-Multi), which are distilled from authentic user issues and replayed within new environments to facilitate rigorous and contamination-free evaluation. Baseline evaluations on BIRD-CRITIC underscore the task’s complexity, with the leading reasoning model O3-Mini achieving only 38.87% success rate on BIRD-CRITIC-PG and 33.33% on BIRD-CRITIC-Multi. Meanwhile, advancing open-source models for database tasks is crucial, empowering local development while safeguarding data privacy. Therefore, we present Six-Gym (Sql-fIX-Gym), a training environment for elevating the capabilities of open-source models specifically for SQL issue debugging. This environment leverages SQL-Rewind strategy, which automatically generates executable issue-solution datasets by reverse-engineering issues from verified SQLs. However, popular trajectory-based fine-tuning methods do not explore substantial supervisory signals. We further propose $f$-Plan Boosting, which extracts high-level debugging plans automatically from SQL solutions, enabling the teacher LLMs to harvest and produce 73.7% more successful trajectories for training. We integrate these components into an open-source agent, Bird-Fixer. Based on Qwen-2.5-Coder-14B, Bird-Fixer raises its success rate to 38.11% on BIRD-CRITIC-PG and 29.65% on BIRD-CRITIC-Multi, surpassing many leading proprietary models such as Claude-3.7-Sonnet and GPT-4.1, marking a significant step toward democratizing sophisticated SQL-debugging capabilities for both research and industry.
@@ 1 Introduction
In this work, we are targeting to bridge this critical gap by two primary contributions. First, we present BIRD-CRITIC, a carefully curated benchmark built from authentic StackOverflow bug-fix threads. It comes in two subsets: (1) BIRD-CRITIC-PG with 530 PostgreSQL-only tasks, and (2) BIRD-CRITIC-Multi, whose 570 tasks are distributed across 4 major open-source dialects containing PostgreSQL, MySQL, SQL Server, and Oracle. Each task undergoes rigorous reconstruction where the underlying knowledge structures and debugging heuristics are extracted, and the scenario is reproduced within a controlled sandbox environment by new RDBs and conditions. This process ensures that tasks remain relevant while minimizing potential exposure to pre-training data. Furthermore, execution accuracy (EX) in standard text-to-SQL is inadequate for the diverse types of issues in BIRD-CRITIC, frequently leading to false negatives. Specifically, tasks involving database state changes, i.e., via Data Manipulation Language (DML) or Data Definition Language operations (DDL), frequently permit multiple functionally equivalent solutions that may differ syntactically or include non-impacting elements [56, 4]; reliance on strict EX matching would incorrectly penalize such valid SQL solutions. Therefore, each task is augmented with custom evaluation scripts containing specific test cases designed to evaluate functional correctness, enabling precise calculation of task success rates. Our baseline evaluations on BIRD-CRITIC underscore the complexity of SQL issue debugging, in which even advanced reasoning models, O3-mini, only achieves a 38.87% success rate on BIRD-CRITIC-PG and 33.33% on BIRD-CRITIC-Multi.
@@ 3 BIRD-CRITIC Benchmark
BIRD-CRITIC is developed via a multi-stage annotation converting raw user issues into executable, verifiable tasks. This involves two annotation groups: 1) 10 qualified database/SQL annotators, who pass strict entry test as detailed in Appendix B.1 and systematic training shown in Appendix B.2 to promise the quality of annotation; 2) 3 senior database experts/scientists for final data collection decision. This process is visually outlined in Figure 2.
@@
We leverage relational databases from the BIRD-SQL development set [26] chosen for its domain diversity across real data-science tasks (California Schools, Financial, Superheroes) and its permissive license. We migrate their original SQLite schemas to PostgreSQL, MySQL, SQL Server, and Oracle, four widely used production-grade dialects. During migration, we go beyond direct dialect translation by refining table and column names. We adjust data types and introduce guarded alterations to schema components to reduce potential information leakage (see Appendix A.2). To pair these databases with realistic debugging scenarios, we collect SQL issue queries from Stack Overflow, following a strict protocol shown in Appendix A.3.
@@
Annotators carefully review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$ can accurately fulfill the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. Also, to ensure robust evaluation, each task is annotated with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix C. We report the Success Rate (SR %), considering a task solved only when $\sigma_{\text{pred}}$ successfully passes all test cases in its evaluation script.
@@
Figure 3: Distribution of issue categories in all BIRD-CRITIC, derived from an analysis of SQL usage in the real-world database applications. A detailed distribution is in Appendix E.
@@
Table 3 summarizes the key properties of the BIRD-CRITIC benchmark, and Figure 3 visualizes the distribution of its underlying knowledge categories. The distribution of benchmark, is detailed in Appendix E. A side-by-side comparison with standard text-to-SQL benchmarks (Table 5 in Appendix E.1) exposes three distinctive challenges introduced by BIRD-CRITIC: non-query-like problems, multi-dialect complexities, and the most verbose but authentic user queries. As far as we know, BIRD-CRITIC is the first debugging benchmark for SQL applications. These aspects establish BIRD-CRITIC as a crucial benchmark for rigorously evaluating LLM proficiency in solving authentic SQL issues.
@@ 4 Six-Gym: An Automated SQL Debugging Environment for LLMs
GYM-like datasets have proven effective for training LLMs as agents for complex tasks [30]. However, manually collecting and annotating these datasets is labor-intensive and difficult to scale, especially for debugging tasks. Thus, we introduce SQL-Rewind, which addresses this by inverting the debugging paradigm: starting with correct SQL queries ($\sigma^{*}$) and systematically introducing realistic issues to generate issue SQLs ($\sigma_{\text{issue}}$) and user issue query $\mathcal{P}$. This approach enables efficient creation of large-scale training data without human annotation. The pseudo-algorithm is shown in Appendix F.1.
@@ 6.1 SetUp
We evaluate the performance of several popular and strong LLMs across two primary categories, including general-purpose models: Gemini-2.0-Flash, GPT-4.1, Claude-3.7-Sonnet, Qwen-2.5-Coder-32B, Meta-Llama-3.1-8B, Meta-Llama-3.3-70B, Phi-4 and DeepSeek-V3. The second category consists of models specifically renowned for their advanced reasoning capabilities: O3-mini, O1-preview, Gemini-2.0-Flash-Thinking, and Claude-3.7-Sonnet-Thinking. The implementation details are in Appendix H.2.
@@ 6.2 Main Results
An analysis of performance across distinct SQL issue categories reveals clear differences in difficulty. Issues related to data management, such as DML operations: insertions, deletions, updates, and DDL operations like schema modifications) are found to be relatively more manageable. On average, reasoning models achieved a 52.6% SR and general-purpose models a 38.8% SR in data management. Issues associated with Personalized functions also demonstrate moderate success rates. In contrast, Query-like issues present the greatest challenge for all LLMs.
@@ 6.5 Error Analysis
To understand how far current LLM-based agents still are from fully resolving user-reported SQL issues, we sample 100 failed tasks from BIRD-CRITIC-PG by 4 agents: O3-mini, GPT-4.1, Claude-3.7-Sonnet, and Bird-Fixer. It can be concluded: Projection Mismatch errors (26.9%), where models misinterpret output requirements by, for instance, adding unexpected columns or misapplying aggregations; Chain of Errors (27.3%), characterized by cascading failures due to partial problem resolution that overlooks dependent issues such as sequence updates accompanying primary key modifications; Incorrect Logic (44.5%), the most prevalent, highlighting fundamental misunderstandings of data structures or transformation methodologies, particularly in complex operations like JSON array manipulation, leading to syntactically plausible but semantically flawed SQL; and Syntax Errors (29.3%), indicating technical implementation flaws such as type mismatches (e.g., DATE versus TIMESTAMP) or improperly formatted intervals, especially in specialized SQL contexts like recursive queries. The detailed examples for each category are in Figure 8.
@@ 8 Conclusion
Appendix A Environment Setup Details
A.1 SQL Dialects Implementation
@@
A.2 Databases Migration & Modifications
@@
A.3 Issue Collection Protocol
@@
Annotators meticulously review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$: (1) Correctly Implements Intent: Accurately fulfills the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. (2) Resolves the Error: Explicitly fixes the identified flaw(s) in $\sigma_{\text{issue}}$. (3) Is Functionally Correct: Executes successfully on the target database instance $D$ (conforming to $\mathcal{S}$) within the specified dialect and produces the expected, correct results. (4) Adheres to Best Practices: Solution SQLs should present a reasonably efficient and well-formed query. As shown in Figure 1, this results in a curated "Solution Query" ($\sigma^{*}$) paired with the user query and issue SQLs. Finally, to ensure robust evaluation, we annotate each task with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix C.
Appendix B Annotator Qualification
B.1 Annotator Entrance Test
@@ B.3 Qualification Test
Appendix C Evaluation Script Details
@@ Appendix D Evaluation Metrics
Appendix E More Statistics
E.1 Comparison of BIRD-CRITIC with other conversational Text-to-SQL benchmarks
@@
E.2 Detailed Statistics of BIRD-CRITIC-Multi
@@
Appendix F Algorithm
F.1 SQL Rewind Algorithm

### v3

@@ Abstract
Resolution of complex SQL issues persists as a significant bottleneck in real-world database applications. Current Large Language Models (LLMs), while adept at text-to-SQL translation, have not been rigorously evaluated on the more challenging task of debugging on SQL issues. In order to address this gap, we introduce BIRD-CRITIC, a new SQL issue debugging benchmark comprising 530 carefully curated PostgreSQL tasks (BIRD-CRITIC-PG) and 570 multi-dialect tasks (BIRD-CRITIC-Multi), which are distilled from authentic user issues and replayed within new environments to facilitate rigorous and contamination-free evaluation. Baseline evaluations on BIRD-CRITIC underscore the task’s complexity, with the leading reasoning model O3-Mini achieving only 38.87% success rate on BIRD-CRITIC-PG and 33.33% on BIRD-CRITIC-Multi. Meanwhile, realizing open-source models for database tasks is crucial which can empower local development while safeguarding data privacy. Therefore, we present Six-Gym (Sql-fIX-Gym), a training environment for elevating the capabilities of open-source models specifically for SQL issue debugging. This environment leverages SQL-Rewind strategy, which automatically generates executable issue-solution datasets by reverse-engineering issues from verified SQLs. However, popular trajectory-based fine-tuning methods do not explore substantial supervisory signals. We further propose $f$-Plan Boosting, which extracts high-level debugging plans automatically from SQL solutions, enabling the teacher LLMs to harvest and produce 73.7% more successful trajectories for training. We integrate these components into an open-source agent, Bird-Fixer. Based on Qwen-2.5-Coder-14B, Bird-Fixer raises its success rate to 38.11% on BIRD-CRITIC-PG and 29.65% on BIRD-CRITIC-Multi, surpassing many leading proprietary models such as Claude-3.7-Sonnet and GPT-4.1, marking a significant step toward democratizing sophisticated SQL-debugging capabilities for both research and industry.
@@ 1 Introduction
In this work, we are targeting to bridge this critical gap by two primary contributions. First, we present BIRD-CRITIC, a carefully curated benchmark built from authentic StackOverflow bug-fix threads. It comes in two subsets: (1) BIRD-CRITIC-PG with 530 PostgreSQL-only tasks, and (2) BIRD-CRITIC-Multi, whose 570 tasks are distributed across 4 major dialects: PostgreSQL and MySQL as open-source databases, SQL Server and Oracle as community-friendly cloud-based platforms with free developer editions. Each task undergoes rigorous reconstruction where the underlying knowledge structures and debugging heuristics are extracted, and the scenario is reproduced within a controlled sandbox environment by new RDBs and conditions. This process ensures that tasks remain relevant while minimizing potential exposure to pre-training data. Furthermore, execution accuracy (EX) in standard text-to-SQL is inadequate for the diverse types of issues in BIRD-CRITIC, frequently leading to false negatives. Specifically, tasks involving database state changes, i.e., via Data Manipulation Language (DML) or Data Definition Language operations (DDL), frequently permit multiple functionally equivalent solutions that may differ syntactically or include non-impacting elements [56, 4]; reliance on strict EX matching would incorrectly penalize such valid SQL solutions. Therefore, each task is augmented with custom evaluation scripts containing specific test cases designed to evaluate functional correctness, enabling precise calculation of task success rates. Our baseline evaluations on BIRD-CRITIC underscore the complexity of SQL issue debugging, in which even advanced reasoning models, O3-mini, only achieves a 38.87% success rate on BIRD-CRITIC-PG and 33.33% on BIRD-CRITIC-Multi.
@@ 3 BIRD-CRITIC Benchmark
BIRD-CRITIC is developed via a multi-stage annotation converting raw user issues into executable, verifiable tasks. This involves two annotation groups: 1) 10 qualified database/SQL annotators, who pass strict entry test as detailed in Appendix 11.1 and systematic training shown in Appendix 11.2 to promise the quality of annotation; 2) 3 senior database experts/scientists for final data collection decision. This process is visually outlined in Figure 2.
@@
We leverage relational databases from the BIRD-SQL development set [26] chosen for its domain diversity across real data-science tasks (California Schools, Financial, Superheroes) and its permissive license. We migrate their original SQLite schemas to PostgreSQL, MySQL, SQL Server, and Oracle, four widely used production-grade dialects. During migration, we go beyond direct dialect translation by refining table and column names. We adjust data types and introduce guarded alterations to schema components to reduce potential information leakage (see Appendix 10.2). To pair these databases with realistic debugging scenarios, we collect SQL issue queries from Stack Overflow, following a strict protocol shown in Appendix 10.3.
@@
Annotators carefully review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$ can accurately fulfill the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. Also, to ensure robust evaluation, each task is annotated with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix 12. We report the Success Rate (SR %), considering a task solved only when $\sigma_{\text{pred}}$ successfully passes all test cases in its evaluation script.
@@
Figure 3: Distribution of issue categories in all BIRD-CRITIC, derived from an analysis of SQL usage in the real-world database applications. A detailed distribution is in Appendix 14.
@@
Table 3 summarizes the key properties of the BIRD-CRITIC benchmark, and Figure 3 visualizes the distribution of its underlying knowledge categories. The distribution of benchmark, is detailed in Appendix 14. A side-by-side comparison with standard text-to-SQL benchmarks (Table 6 in Appendix 14.1) exposes three distinctive challenges introduced by BIRD-CRITIC: non-query-like problems, multi-dialect complexities, and the most verbose but authentic user queries. As far as we know, BIRD-CRITIC is the first debugging benchmark for SQL applications. These aspects establish BIRD-CRITIC as a crucial benchmark for rigorously evaluating LLM proficiency in solving authentic SQL issues.
@@ 4 Six-Gym: An Automated SQL Debugging Environment for LLMs
GYM-like datasets have proven effective for training LLMs as agents for complex tasks [30]. However, manually collecting and annotating these datasets is labor-intensive and difficult to scale, especially for debugging tasks. Thus, we introduce SQL-Rewind, which addresses this by inverting the debugging paradigm: starting with correct SQL queries ($\sigma^{*}$) and systematically introducing realistic issues to generate issue SQLs ($\sigma_{\text{issue}}$) and user issue query $\mathcal{P}$. This approach enables efficient creation of large-scale training data without human annotation. The pseudo-algorithm is shown in Appendix 17.1.
@@ 6.1 SetUp
We evaluate the performance of several popular and strong LLMs across two primary categories, including general-purpose models: Gemini-2.0-Flash, GPT-4.1, Claude-3.7-Sonnet, Qwen-2.5-Coder-32B, Meta-Llama-3.1-8B, Meta-Llama-3.3-70B, Phi-4 and DeepSeek-V3. The second category consists of models specifically renowned for their advanced reasoning capabilities: O3-mini, O1-preview, Gemini-2.0-Flash-Thinking, and Claude-3.7-Sonnet-Thinking. The implementation details are in Appendix 16.2.
@@ 6.2 Main Results
An analysis of performance across distinct SQL issue categories reveals clear differences in difficulty. Issues related to data management, such as DML operations: insertions, deletions, updates, and DDL operations like schema modifications, are found to be relatively more manageable. On average, reasoning models achieved a 52.6% SR and general-purpose models a 38.8% SR in data management. Issues associated with Personalized functions also demonstrate moderate success rates. In contrast, Query-like issues present the greatest challenge for all LLMs.
@@ 6.3 Performance Analysis of Bird-Fixer
6.4 Trajectory Sampling Comparison
Table 4: Trajectory Generation Efficiency Comparison. Baseline: Standard SQL-ACT rollout with a single attempt (temperature=0). f-Plan (Ours): A single rollout guided by functional plans extracted from issue–solution pairs (temperature=0). Rejection Sampling: Up to 5 trials per instance (temperature=0.8), with early stopping when a successful trajectory is obtained. Reject + f-Plan: Combination of rejection sampling (up to 5 trials) with f-Plan guidance.
Method Max Successful Avg DB Time Cost
Tries Traj. Tries (min) (\$)
Baseline 1 1,254 1.0 306 8.47
f-Plan 1 2,178 1.0 324 27.44
Rejection Sampling 5 1,910 4.2 1,377 108.05
Reject + f-Plan 5 2,560 1.7 810 41.16
To better illustrate the efficiency and effectiveness of f-Plan Boosting, we compare it against widely used trajectory augmentation approaches. We evaluate four strategies for trajectory generation, using Gemini-2.0-Flash as the teacher model on Six-Gym.
As shown in Table 4, f-Plan Boosting yields $73.7\%$ more successful trajectories than the baseline while maintaining similar runtime and overhead. By contrast, rejection sampling increases success rates modestly but at the cost of $4.2\times$ more attempts and $4.5\times$ longer execution time. When combined, rejection sampling and f-Plan achieve the best overall trade-off, generating 2,560 trajectories with reduced average attempts ($1.7$) and a $62\%$ reduction in cost relative to rejection sampling alone. These results demonstrate that f-Plan provides an effective and efficient approach to trajectory augmentation during rollout in complex environments. Other detailed comparison can be found in Appendix 14.3.
@@ 6.6 Error Analysis
To understand how far current LLM-based agents still are from fully resolving user-reported SQL issues, we sample 100 failed tasks from BIRD-CRITIC-PG by 4 agents based on: O3-mini, GPT-4.1, Claude-3.7-Sonnet, and Bird-Fixer. It can be concluded that current agents exhibit four distinct error modes reflecting different levels of reasoning deficiency: Projection Mismatch errors (26.9%), where models misinterpret output requirements by, for instance, adding unexpected columns or misapplying aggregations, suggesting limitations in semantic understanding of user intent and schema alignment; Chain of Errors (27.3%), characterized by cascading failures due to partial problem resolution that overlooks dependent issues such as sequence updates accompanying primary key modifications, revealing difficulties in multi-step causal reasoning and consistency maintenance; The database engine only reports the most superficial issue, masking a deeper, dependent error that is the true root cause. For instance, a type mismatch error might be reported, but the underlying problem could be an incorrect join that brought together the wrong columns in the first place. Incorrect Logic (44.5%), the most prevalent, highlighting fundamental misunderstandings of data structures or transformation methodologies, particularly in complex operations like JSON array manipulation, leading to syntactically plausible but semantically flawed SQL; and Syntax Errors (29.3%), indicating technical implementation flaws such as type mismatches (e.g., DATE versus TIMESTAMP) or improperly formatted intervals, especially in specialized SQL contexts like recursive queries. The detailed examples for each category are in Figure 8. These findings highlight that future improvements should emphasize logical and schema-aware reasoning, cross-step dependency tracking, and dialect-robust SQL generation rather than mere syntactic refinement.
@@ 8 Conclusion
9 Acknowledgments
We thank the anonymous reviewers and committees for their helpful comments, suggestions and organizations. We thank John Yang for early discussion and suggestions. Reynold Cheng, Jinyang Li, Xiaolong Li, Ge Qu, Nan Huo, Xiaohan Xu, and Ziwei Tang were supported by the Research Grant Council of Hong Kong (RGC Project HKU 17202325), the University of Hong Kong (Project 2409100399), and the HKU Faculty Exchange Award 2024 (Faculty of Engineering).
Appendix Contents
10 Environment Setup Details
10.1 SQL Dialects Implementation
@@
10.2 Databases Migration & Modifications
@@
10.3 Issue Collection Protocol
@@
Annotators meticulously review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$: (1) Correctly Implements Intent: Accurately fulfills the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. (2) Resolves the Error: Explicitly fixes the identified flaw(s) in $\sigma_{\text{issue}}$. (3) Is Functionally Correct: Executes successfully on the target database instance $D$ (conforming to $\mathcal{S}$) within the specified dialect and produces the expected, correct results. (4) Adheres to Best Practices: Solution SQLs should present a reasonably efficient and well-formed query. As shown in Figure 1, this results in a curated "Solution Query" ($\sigma^{*}$) paired with the user query and issue SQLs. Finally, to ensure robust evaluation, we annotate each task with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix 12.
11 Annotator Qualification
11.1 Annotator Entrance Test
@@ 11.3 Qualification Test
12 Evaluation Script Details
@@ 13 Evaluation Metrics
14 More Statistics
14.1 Comparison of BIRD-CRITIC with other conversational Text-to-SQL benchmarks
@@
14.2 Detailed Statistics of BIRD-CRITIC-Multi
@@
14.3 Quality Validation of Six-Gym
This section validates the quality of synthetic data generated by SQL-Rewind by comparing Six-Gym with the manually curated BIRD-CRITIC-PG benchmark. Table 8 demonstrates that our synthetic dataset exhibits comparable complexity and diversity across multiple dimensions, including similar distribution of complex operations, higher SQL diversity ratio, and comparable query lengths to human-annoted challenging data.
Table 8: Data Statistics Comparison between Six-Gym and BIRD-CRITIC-PG $\dagger$}]: Diversity Ratio = Unique 3-grams / Total 3-grams.
Dimension BIRD-CRITIC-PG SIX-GYM
User Query Length (mean/max) 162.98/1046 171.1/882
Issue SQL Length (mean/max) 133.29/1262 110.2/1089
Solution SQL Length (mean/max) 112.64/853 94.8/772
SQL Keywords Coverage 165 157
Complex Operations (%) 54.5 54.3
Multi-clause Queries (%) 59.4 61.2
SQL Diversity Ratio† 0.728 0.750
Table 9 further breaks down performance by SQL complexity. The benefits of f-Plan scale with difficulty: while gains over rejection sampling are modest on simple queries (+7.5 points), they grow dramatically on complex queries with $5+$ clauses (+29.2 points). f-Plan also resolves instances unsolved by either baseline or rejection sampling, particularly those with high keyword diversity and nested operations. These results highlight that f-Plan narrows the search space through structured debugging plans, providing explicit guidance that is especially valuable when random exploration becomes ineffective.
Table 9: Success Rate by SQL Complexity
Issue SQL Complexity Baseline Reject f-Plan
Simple (1-2 clauses) 52.3% 62.8% 70.3%
Medium (3-4 clauses) 38.7% 58.8% 69.4%
Complex (5+ clauses) 19.4% 37.1% 66.3%
High Keyword Diversity ($10+$) 24.1% 35.6% 54.3%
Nested Operations ($2+$ levels) 21.8% 36.8% 49.2%
15 Error Analysis Details
Figure 8: Detailed Error Analysis
16 Experiment Details
16.1 Alias of LLMs
The following aliases are used for the models in this work:
  • Claude-3.7-Sonnet: claude-3-7-sonnet-20250219
  • O3-Mini: O3-Mini-2025-01-31
  • O1-Preview: O1-Preview-2024-09-12
  • GPT-4.1: gpt-4.1-2025-04-14
  • Gemini-2.0-Flash: gemini-2.0-flash
  • deepseek-v3: deepseek-chat
  • deepseek-r1: deepseek-reasoner
  • Phi: Phi-4
16.2 Model Implementation Details
[…]
Table 10: GPU hours spent to train each backbone model.
Model GPU Hours
Meta-Llama-3.1-8B 24.88
Qwen2.5-Coder-7B 22.00
Qwen2.5-Coder-14B 35.93
Phi-4 31.42
16.3 Agent Implementation Details
  • Tool-ACT: Actions are predefined and include:
    • Schema Inspection: Reveals table/column information.
    • Sample Data: Previews example rows from a table.
17 Algorithm
17.1 SQL Rewind Algorithm